OpenStudy (anonymous):

Forty-five is 20% of what number? help

OpenStudy (anonymous):

thanks for helping

OpenStudy (graciebugg):

Its 225

OpenStudy (anonymous):

how do you know???

OpenStudy (anonymous):

to get the percent of a number, you can do: # * 0.2 = 45 so, to go backwards, you can do 45 / 0.2 = #

OpenStudy (anonymous):

ok so who is right you of the other pirsion

OpenStudy (anonymous):

@GracieBugg you're not supposed to give direct answers !

OpenStudy (alexandervonhumboldt2):

x*25/100=45

OpenStudy (anonymous):

ok

OpenStudy (alexandervonhumboldt2):

x=180

OpenStudy (anonymous):

i have 4 more questions

OpenStudy (alexandervonhumboldt2):

ok

OpenStudy (anonymous):

180 is not correct. 225 is.

OpenStudy (anonymous):

???

OpenStudy (alexandervonhumboldt2):

woops i mistaked 25 with 20%

OpenStudy (alexandervonhumboldt2):

x*20/100=45

OpenStudy (anonymous):

it's okay (:

OpenStudy (anonymous):

ok next one is Mai bought a television for $325. The tax rate is 8%. Which proportion can be used to find the amount of tax Mai will pay? A. B. C. D.

OpenStudy (graciebugg):

I didn't know the direct answer thing

OpenStudy (anonymous):

it's okay !! the answer choices aren't there.